Martin Company currently produces and sells 41,000 units of product at a selling price of $15. The product has variable costs of $8 per unit and fixed costs of $51,000. The company currently earns a total contribution margin of:

Multiple Choice

$357,000

$459,000

$287,000

$408,000
Answer #1

contribution margin=Selling price-Variable cost

=(15-8)=$7 per unit

Total contribution margin=(7*41000)

=$287,000